OpenStudy (anonymous):

MEDAL + FAN Simplify (-1 + i)^4.

OpenStudy (studygurl14):

\(\large (a+b)^2 = a^2+2ab+b^2\) So... \(\large (a+b)^4=(a+b)^2(a+b)^2 = (a^2+2ab+b^2)(a^2+2ab+b^2)\) In this situation, a = -1, and b=i Plug in and simplify. Let me know if you run into any trouble. :)